# From Canopy to Ground via ForestGen3D: Learning Cross-Domain Generation of 3D Forest Structure from Aerial-to-Terrestrial LiDAR

### v1

@@ Abstract
The 3D structure of living and non-living components in ecosystems plays a critical role in determining ecological processes and feedbacks from both natural and human-driven disturbances. Anticipating the effects of wildfire, drought, disease, or atmospheric deposition depends on accurate characterization of 3D vegetation structure, yet widespread measurement remains prohibitively expensive and often infeasible. We introduce ForestGen3D, a novel generative modeling framework that synthesizes high-fidelity 3D forest structure using only aerial LiDAR (ALS) inputs. ForestGen3D is based on conditional denoising diffusion probabilistic models (DDPMs) trained on co-registered ALS/TLS (terrestrial LiDAR) data. The model learns to generate TLS-like 3D point clouds conditioned on sparse ALS observations, effectively reconstructing occluded sub-canopy detail at scale. To ensure ecological plausibility, we introduce a geometric containment prior based on the convex hull of ALS observations and provide theoretical and empirical guarantees that generated structures remain spatially consistent. We evaluate ForestGen3D at tree, plot, and landscape scales using real-world data from mixed conifer ecosystems, and show that it produces high-fidelity reconstructions that closely match TLS references in terms of geometric similarity and biophysical metrics, such as tree height, DBH, crown diameter and crown volume. Additionally, we demonstrate that the containment property can serve as a practical proxy for generation quality in settings where TLS ground truth is unavailable. Our results position ForestGen3D as a scalable tool for ecological modeling, wildfire simulation, and structural fuel characterization in ALS-only environments.
@@ 1 Introduction
Wildfires, drought, pathogens, and other natural disturbances are increasingly transforming forested landscapes, driving the need for improved ecological monitoring and fuel structure characterization. These disturbances are shaped not only by vegetation type but also by the three-dimensional (3D) arrangement of biomass across vertical and horizontal layers. The 3D structure of both living and non-living components of vegetation plays a critical role in regulating ecological processes, including energy and moisture fluxes, carbon storage, and fire behavior.
@@
However, large-scale collection of detailed 3D vegetation structure remains a major bottleneck. Most landscape-scale ecological assessments still rely on field plots that are statistically upscaled using summary metrics. To spatially represent forests, synthetic layouts are often generated by populating plots with idealized geometric models (e.g., cylinders, ovals), randomly placed according to field statistics [1, 2, 3, 4]. While informative, these methods often overlook spatial relationships between vegetation components and fail to capture the full 3D spatial heterogeneity, both of which are crucial for modeling dynamic ecological processes. A related class of methods common in visualization, and gaming applications uses procedural forest generation, where tree forms and vegetation distributions are instantiated from parametric rules or stylized templates. Tools such as SpeedTree [5] and Unity’s Terrain Engine [6] exemplify this approach, offering efficient generation of stylized vegetation for rendering purposes. While these models can quickly populate large virtual landscapes, they typically prioritize visual coherence and rendering performance over fidelity with real-world forest conditions, and are therefore less suitable for scientific analyses that depend on detailed and ecologically grounded 3D representations of forest fuel structure.
@@
In this work, we propose to bridge these sensing gaps by learning generative models of 3D fuel structure that leverage the strengths of both data sources. To achieve this, we develop ForestGen3D, a cross-domain generative modeling framework based on denoising diffusion probabilistic models (DDPMs). These models are trained here on a new dataset of co-registered ALS/TLS tree 3D point clouds referred to as the CoLiDAR Forest3D (ALS/TLS) and collected across mixed conifer vegetation ecosystems. Once trained, the model can generate novel 3D vegetation samples by starting from random noise and iteratively refining structure through a learned denoising process, guided by ALS input as shown in Fig. 2. To the best of our knowledge, this is the first approach to leverage conditional diffusion models for cross-domain generation of 3D forest structure, bridging the gap between sparse, landscape-scale ALS and detailed, plot-scale, TLS representations. This enables scalable ecological modeling of forests in full vertical profile beyond existing upsampling or procedural simulation methods.
@@
Figure 2: Denoising Diffusion Probabilistic Model (DDPM) in Action. DDPM iteratively denoises a sample guided by ALS measurements and generates a sample of forest tree 3D structure from the learned training distribution. Note that the 3D structural information provided by ALS is effective for constraining 3D structural generation in the vertical tree stand direction.
@@
Existing 3D generative and completion models, such as PCN [14], 3D-GAN [15], latent-GAN [16], and PointFlow [17], have so far been developed and tested primarily on man-made objects from synthetic datasets such as ShapeNet. While effective in those controlled domains, their suitability for modeling real-world forest vegetation remains largely unknown. Forest structure presents fundamentally different challenges: complex topologies with intertwined crowns, occlusion-driven sparsity due to canopy cover, and vertical heterogeneity across understory, stems, and crowns. These characteristics may degrade the performance of models designed for compact, uniformly sampled shapes. Moreover, existing methods are typically trained without reference to real-world sensing modalities such as LiDAR, and therefore do not account for the anisotropic density, partial observability, and sensor-specific artifacts present in ecological data. In contrast, the objective of our approach is to generate ecologically realistic 3D vegetation structure directly from ALS/TLS data collected in mixed conifer forests, capturing both the natural variability of vegetation and the sensor-dependent effects inherent in terrestrial and airborne LiDAR measurements.
In the sections that follow, we present our dataset and training approach, describe the DDPM-based model, and evaluate ForestGen3D across tree, plot, and landscape scale experiments. Quantitative and qualitative results demonstrate that our approach approximates TLS-level structural detail using ALS data alone, enabling scalable forest modeling for applications in fire science, ecological monitoring, and remote sensing.
@@ 2 ForestGen3D
The core problem we address is generating synthetic forests by improving the fidelity of individual tree structures beyond conventional coarse-shape models. To achieve this, we explicitly preserve ecosystem fidelity by conditioning the generation process on ALS inputs, which provide canopy-level structure tied to real forest layouts. Rather than modeling entire plots or landscapes directly which can compromise fine-grained detail and impose significant computational demands, we focus on learning 3D structure at the tree scale, where the physical representation of vegetation is most accurate. This localized modeling approach allows for more efficient training and data preparation, reducing the need to capture broad ecosystem variation in a single model. Moreover, it enables scalable deployment by applying the generative model independently to each detected tree at the plot or landscape scale, under the assumption that only regions beneath vegetation are susceptible to occlusion in ALS data. To support this effort, we constructed a 3D forest dataset of co-registered ALS/TLS tree samples, which we used to train cross-domain denoising diffusion probabilistic models (DDPMs). These models learn the statistical distribution of tree 3D structure in a target domain (TLS) conditioned on observations in a reference domain (ALS). DDPMs start from random Gaussian noise and iteratively refine structure through a learned denoising process, enabling flexible and efficient generation of high-resolution 3D tree forms. Fundamentally, the framework learns to generate samples from the TLS distribution conditioned on a measured sample from the ALS distribution. Figure 2 illustrates this process: an ALS input guides the denoising steps, which progressively reconstruct the full 3D structure of a tree, including stem and understory geometry missing from the input.
@@ 2.1 Training dataset
We construct a new 3D forest dataset CoLiDAR-Forest3D (ALS/TLS) dataset, containing 2,900 co-registered TLS/ALS tree scans for training, including both individual trees and clusters of closely packed trees. The dataset was constructed using the co-registration approach from [18] to align TLS scans with overlapping ALS data and the deep-learning-based tree detection method from [19] to extract individual trees. The raw data were collected from a mixed-conifer forest at Fort Stewart in the United States [20], primarily featuring species such as longleaf pine, slash pine, loblolly pine, and turkey oak. The understory consists of graminoids, forbs, and shrubs, including wiregrass, gallberry, and saw palmetto.
@@
Figure 3: 9 examples of the 3D training dataset consisting of 2900 examples of segmented and co-registered ALS/TLS trees. ALS trees in yellow in the top row, corresponding co-registered TLS trees in aqua color in the bottom row.
@@
TLS data was acquired using a Leica BLK360 LiDAR scanner with an 830 nm wavelength laser and a field of view (FOV) of 300° (vertical) and 360° (horizontal). Scans were taken from static tripod positions, with 240 scans randomly distributed across a 2.5 km diameter area. Each TLS scan contained approximately 8 million points, with around 4.6 million from the ground. Since most returns were concentrated within a 25-meter range, points beyond this threshold were removed to mitigate the non-uniform point density. A representative sampling pattern of TLS scans across a region is shown in Fig. 1, where each TLS scan is represented by a color-coded blob. ALS data was collected using a Galaxy T2000 LiDAR sensor mounted on a fixed-wing aircraft. The sensor operates at a 1064 nm wavelength and captures up to eight returns per pulse, with a 28° FOV, flown at an altitude of approximately 1,675 meters. The system achieves a point density of approximately 15 points per square meter and maintains an accuracy of 0.03 to 0.25 meters RMSE within a 150–6,500 meter range. The surveyed area was covered using multiple overlapping flight lines, including seven cross strips to reduce point sparsity and enhance coverage. Unlike TLS, ALS points are not affected by satellite signal occlusion and were globally aligned using a Trimble CenterPoint RTX RTK system for centimeter-level accuracy. A regional-scale ALS point cloud is shown in Fig. 1 in yellow, with orthographic (top-down) views on the left and middle and a side-view on the top right.
@@
Each TLS scan was initially positioned using GPS-based estimates in global coordinates. These positions were used to extract a corresponding 50 meter radius ALS point-cloud subset around each TLS scan center. The automatic and target-less co-registration approach from [18] was applied to refine the position provided by GPS while also aligning the TLS scan to the ALS point-cloud relative to rotations. Each co-registered pair was manually verified by an expert to ensure qualitative alignment between spatial features. Tree detection was performed using the deep learning (DL)-based method from [19], which detects trees or tightly clustered groups of trees/vegetation by fitting 2D bounding boxes from a top-down bird’s-eye view and then expanding each bounding box to 3D informed by the lowest-to-highest point elevation difference. This approach ensures that the resulting 3D bounding boxes capture the full vertical extent of vegetation, from the top of the canopy to the lowest ground point. These bounding boxes were used to extract tree-level point clouds by truncating the points in the ALS and TLS scans inside each 3D box. To mitigate the non-uniform point density in TLS, we truncated each point cloud to a 25m radius and downsampled each TLS scan to a 0.10 cm voxel size. Each tree scan was then centered and normalized within the $0,1]^{3}$ range. The dataset was manually curated to remove incomplete trees (e.g., those significantly cut by bounding boxes), and incorrect detections. The final dataset consists of 2,900 TLS/ALS tree or tree-cluster pairs. Fig. 3 illustrates nine ALS/TLS example pairs, with ALS trees in yellow (top row) and corresponding TLS scans in aqua (bottom row). For validation, we use an additional set of 300 tree samples from the same ecosystem than the training data but that are non-overlapping. These examples were collected using the same TLS and ALS sensors and processed using identical co-registration and tree detection pipelines.
Figure 4: TLS containment within ALS convex envelope
@@
for a small constant $\epsilon\ll 1$, with $(\mathbf{X},\mathbf{Y})\sim\mathcal{D}$ drawn from the dataset of co-registered tree-level TLS/ALS pairs. Here, $\mathbf{x}\sim\mathbf{X}$ are 3D points uniformly sampled from the TLS point cloud, and $1(\cdot)$ is the indicator function. In our empirical analysis of the training CoLIDAR-Forest3D (ALS/TLS) dataset, approximately 3.9% of TLS points fall outside the ALS convex hull, i.e., $\epsilon=0.04$, validating the assumption at the tree scale. We emphasize that this prior is not assumed to hold globally at the plot or landscape scale and that a variety of factors may violate containment. However, the small deviations from containment that may be locally present are expected and well-justified, as they can result from sensor misalignment, resolution differences, vegetation visible from the ground but occluded from the aerial view, and inherent measurement noise in both ALS and TLS systems. Despite these localized discrepancies, the overall containment expectation introduces a geometric prior that is implicitly learned by our conditional generative model. This prior not only helps constrain synthesis within plausible ALS-guided structure, but also provides a useful means of measuring the spatial fidelity of the generated points relative to the ALS envelope.
@@
The DDPM architecture consists of two components: a forward diffusion process that progressively corrupts the input point cloud $\mathbf{X}^{0}\in\mathbb{R}^{N\times 3}$ over $T$ steps by multiplying it with a decaying factor while adding Gaussian noise, and a learned reverse process that denoises the corrupted sample back to a clean reconstruction. This yields a training trajectory: $\mathbf{X}^{0}\rightarrow\mathbf{X}^{1}\rightarrow\cdots\rightarrow\mathbf{X}^{T}\sim\mathcal{N}(0,\mathbf{I})$, as the forward (diffusion) process, $\mathbf{X}^{T}\rightarrow\mathbf{X}^{T-1}\rightarrow\cdots\rightarrow\hat{\mathbf{X}}^{0}$ as the reverse (denoising) process. The model conditions each denoising step on ALS observations $\mathbf{Y}$, enabling it to learn structure consistent with the canopy geometry while generating plausible TLS content beneath. The network uses an encoder–decoder U-Net structure adapted to point clouds [22], where each block is implemented using PointNet++ layers [23]. Conditioning is applied through feature-wise modulation blocks at each level of the hierarchy, allowing ALS input to guide the generation process. A high-level schematic of the model architecture is shown in Fig. 5.
Figure 5: Schematic of ALS-to-TLS cross-domain DDPM. The forward process encodes a TLS sample into Gaussian noise, while the reverse process learns to denoise a sample conditioned on ALS input, reconstructing high-resolution 3D vegetation structure.

### v2

@@ Abstract
The 3D structure of living and non-living components in ecosystems plays a critical role in determining ecological processes and feedbacks from both natural and human-driven disturbances. Anticipating the effects of wildfire, drought, disease, or atmospheric deposition depends on accurate characterization of 3D vegetation structure, yet widespread measurement remains prohibitively expensive and often infeasible. We present ForestGen3D, a cross-domain generative framework that preserves ALS observed 3D forest structure while inferring missing sub-canopy detail. ForestGen3D is based on conditional denoising diffusion probabilistic models trained on co-registered ALS and terrestrial LiDAR (TLS) data, exploiting the empirical containment of TLS structure within the ALS derived spatial envelope. The model generates realistic TLS-like point clouds that remain spatially consistent with ALS geometry, enabling landscape-scalable reconstruction of full vertical forest structure. To ensure ecological plausibility, we introduce a geometric containment prior based on the convex hull of ALS observations and provide theoretical and empirical guarantees that generated structures remain spatially consistent. We evaluate ForestGen3D at tree, plot, and landscape scales using real-world data from mixed conifer ecosystems, and show through qualitative and quantitative geometric and distributional analyses that it produces high-fidelity reconstructions closely matching TLS reference data in terms of 3D structural similarity and downstream biophysical metrics, including tree height, DBH, crown diameter, and crown volume. We further demonstrate that the resulting expected point containment (EPC) serves as a practical proxy for generation quality in settings where TLS ground truth is unavailable. Our results demonstrate that ForestGen3D enhances the utility of ALS only environments by inferring ecologically plausible sub-canopy structure while faithfully preserving the landscape heterogeneity encoded in ALS observations, thereby providing a richer 3D representation for ecological analysis, structural fuel characterization and related remote sensing applications.
@@ 1 Introduction
Wildfires, drought, pathogens, and other natural disturbances are increasingly transforming forested landscapes, driving the need for improved ecological monitoring and fuel structure characterization. These disturbances are shaped not only by vegetation type but also by the three-dimensional (3D) arrangement of biomass across vertical and horizontal layers. The 3D structure of both living and non-living components of vegetation plays a critical role in regulating ecological processes, including energy and moisture fluxes, carbon storage, and fire behavior [1, 2, 3, 4].
@@
However, large-scale collection of detailed 3D vegetation structure remains a major bottleneck. Most landscape-scale ecological assessments still rely on field plots that are statistically upscaled using summary metrics [5]. To spatially represent forests, synthetic layouts are often generated by populating plots with idealized geometric models (e.g., cylinders, ovals), randomly placed according to field statistics [6, 7, 8, 9]. While informative, these methods often overlook spatial relationships between vegetation components and fail to capture the full 3D spatial heterogeneity, both of which are crucial for modeling dynamic ecological processes. A related class of methods common in visualization, and gaming applications uses procedural forest generation, where tree forms and vegetation distributions are instantiated from parametric rules or stylized templates. Tools such as SpeedTree [10] and Unity’s Terrain Engine [11] exemplify this approach, offering efficient generation of stylized vegetation for rendering purposes. While these models can quickly populate large virtual landscapes, they typically prioritize visual coherence and rendering performance over fidelity with real-world forest conditions, and are therefore less suitable for scientific analyses that depend on detailed and ecologically grounded 3D representations of forest fuel structure.
@@
In this work, we propose to bridge these sensing gaps by learning generative models of 3D fuel structure that leverage the strengths of both data sources. To achieve this, we develop ForestGen3D, a cross-domain generative modeling framework based on denoising diffusion probabilistic models (DDPMs). DDPMs are a class of deep generative models that learn the underlying data distribution by gradually adding noise to training examples and then learning to reverse this process through iterative denoising [19, 20]. Once trained, the model can synthesize new samples by starting from pure Gaussian noise and progressively refining structure in a manner consistent with the learned distribution. In our case, the DDPM is conditioned on ALS inputs so that the denoising process respects the geometric information encoded in airborne LiDAR while generating plausible sub-canopy structure. These models are trained here on a new dataset of co-registered ALS/TLS tree 3D point clouds referred to as the CoLiDAR Forest3D (ALS/TLS) and collected across mixed conifer vegetation ecosystems. Once trained, the model can generate novel 3D vegetation samples by starting from random noise and iteratively refining structure through a learned denoising process, guided by ALS input as shown in Fig. 2. To the best of our knowledge, this is the first approach to leverage conditional diffusion models for cross-domain generation of 3D forest structure, bridging the gap between sparse, landscape-scale ALS and detailed, plot-scale, TLS representations. This enables scalable ecological modeling of forests in full vertical profile beyond existing upsampling or procedural simulation methods.
@@
Figure 2: ForestGen3D based on denoising diffusion probabilistic model (DDPM) formulation. Model iteratively denoises a sample guided by ALS measurements and generates a sample of forest tree 3D structure from the learned training distribution. Note that the 3D structural information provided by ALS is effective for constraining 3D structural generation in the vertical tree stand direction.
@@
Existing 3D generative and completion models, such as PCN [21], 3D-GAN [22], latent-GAN [23], and PointFlow [24], have so far been developed and tested primarily on man-made objects from synthetic datasets such as ShapeNet. While effective in those controlled domains, their suitability for modeling real-world forest vegetation remains largely unknown. Forest structure presents fundamentally different challenges: complex topologies with intertwined crowns, occlusion-driven sparsity due to canopy cover, and vertical heterogeneity across understory, stems, and crowns. These characteristics may degrade the performance of models designed for compact, uniformly sampled shapes. Moreover, existing methods are typically trained without reference to real-world sensing modalities such as LiDAR, and therefore do not account for the anisotropic density, partial observability, and sensor-specific artifacts present in ecological data. In contrast, the objective of our approach is to generate ecologically realistic 3D vegetation structure directly from ALS/TLS data collected in mixed conifer forests, capturing both the natural variability of vegetation and the sensor dependent effects inherent in terrestrial and airborne LiDAR measurements. In the sections that follow, we present our dataset and training approach, describe the DDPM-based model, and evaluate ForestGen3D across tree, plot, and landscape scale experiments. Quantitative and qualitative results demonstrate that the proposed approach approximates TLS-level structural detail using ALS data alone.
@@ 2 ForestGen3D
The core problem we address is generating synthetic forests by improving the fidelity of individual tree structures beyond conventional coarse-shape models. To achieve this, we explicitly preserve ecosystem fidelity by conditioning the generation process on ALS inputs, which provide canopy level structure tied to real forest layouts. Rather than modeling entire plots or landscapes directly which can compromise fine-grained detail and impose significant computational demands, we focus on learning 3D structure at the tree scale, where the physical representation of vegetation is most accurate. This localized modeling approach allows for more efficient training and data preparation, reducing the need to capture broad ecosystem variation in a single model. Moreover, it enables scalable deployment by applying the generative model independently to each detected tree at the plot or landscape scale, under the assumption that only regions beneath vegetation are susceptible to occlusion in ALS data. To support this effort, we constructed a 3D forest dataset of co-registered ALS/TLS tree samples, which we used to train cross-domain denoising diffusion probabilistic models (DDPMs). These models learn the statistical distribution of tree 3D structure in a target domain (e.g., TLS) conditioned on observations in a reference domain (e.g., ALS). DDPMs start from random Gaussian noise and iteratively refine structure through a learned denoising process, enabling flexible and efficient generation of high-resolution 3D tree forms. Fundamentally, the framework learns to generate samples from the TLS distribution conditioned on a measured sample from the ALS distribution. Figure 2 illustrates this process: an ALS input guides the denoising steps, which progressively reconstruct the full 3D structure of a tree, including stem and understory geometry missing from the input.
@@ 2.1 Training dataset
We construct a new 3D forest dataset CoLiDAR-Forest3D (ALS/TLS) dataset, containing 2,900 co-registered TLS/ALS tree scans for training, including both individual trees and clusters of closely packed trees. The dataset was constructed using the co-registration approach from [25] to align TLS scans with overlapping ALS data and the deep-learning based tree detection method from [26] to extract individual trees. The raw data were collected from a mixed-conifer forest at Fort Stewart in the United States [27], primarily featuring species such as longleaf pine, slash pine, loblolly pine, and turkey oak. The understory consists of graminoids, forbs, and shrubs, including wiregrass, gallberry, and saw palmetto.
@@
Figure 3: 9 examples of the 3D training dataset consisting of 2900 examples of tree-segmented and co-registered aerial lidar scanning (ALS)/ terrestrial lidar scanning (TLS) trees. ALS trees in yellow in the top row, corresponding co-registered TLS trees in aqua color in the bottom row.
@@
TLS data was acquired using a Leica BLK360 LiDAR scanner with an 830 nm wavelength laser and a field of view (FOV) of 300° (vertical) and 360° (horizontal). Scans were collected using a single-scan acquisition strategy, in which the instrument is deployed once at a plot centered position rather than through multiple co-registered scans. Although single viewpoint TLS does not fully eliminate occlusion, especially in dense forest understories, this strategy has been shown to be both effective and efficient for characterizing tree and plot level structural attributes across a wide range of forest ecosystems. Prior work demonstrates that randomized single-scan deployments provide reliable summary estimates of forest characterizations, while enabling practical data collection at larger spatial and ecological extents [12, 28, 29]. Consequently, our dataset prioritizes broader ecological coverage and diversity of vegetation conditions rather than full multi-view TLS completeness, aligning with the objective of training a generative model that generalizes across landscape heterogeneity. Scans were taken from static tripod positions, with 240 scans randomly distributed across a 2.5 km diameter area. Each TLS scan contained approximately 8 million points, with around 4.6 million from the ground. Since most returns were concentrated within a 25 meter range, points beyond this threshold were removed to mitigate the non-uniform point density. A representative sampling pattern of TLS scans across a region is shown in Fig. 1, where each TLS scan is represented by a color-coded blob. ALS data was collected using a Galaxy T2000 LiDAR sensor mounted on a fixed-wing aircraft. The sensor operates at a 1064 nm wavelength and captures up to eight returns per pulse, with a 28° FOV, flown at an altitude of approximately 1,675 meters. The system achieves a point density of approximately 15 points per square meter and maintains an accuracy of 0.03 to 0.25 meters RMSE within a 150–6,500 meter range. The surveyed area was covered using multiple overlapping flight lines, including seven cross strips to reduce point sparsity and enhance coverage. Unlike TLS, ALS points are not affected by satellite signal occlusion and were globally aligned using a Trimble CenterPoint RTX RTK system for centimeter-level accuracy. A regional scale ALS point cloud is shown in Fig. 1 in yellow, with orthographic (top-down) views on the left and middle and a side-view on the top right.
@@
Each TLS scan was initially positioned using GPS-based estimates in global coordinates. These positions were used to extract a corresponding 50 meter radius ALS pointcloud subset around each TLS scan center. The automatic and targetless co-registration approach from [25] was applied to refine the position provided by GPS while also aligning the TLS scan to the ALS point-cloud relative to rotations. Each co-registered pair was manually verified by an expert to ensure qualitative alignment between spatial features. Tree detection was performed using the deep learning (DL) based method from [26], which detects trees or tightly clustered groups of trees/vegetation by fitting 2D bounding boxes from a top-down bird’s-eye view and then expanding each bounding box to 3D informed by the lowest-to-highest point elevation difference. This approach ensures that the resulting 3D bounding boxes capture the full vertical extent of vegetation, from the top of the canopy to the lowest ground point. These bounding boxes were used to extract tree-level point clouds by truncating the points in the ALS and TLS scans inside each 3D box. To mitigate the non-uniform point density in TLS, we truncated each point cloud to a 25m radius and downsampled each TLS scan to a 0.10 cm voxel size. Each tree scan was then centered and normalized within the $0,1]^{3}$ range. The dataset was manually curated to remove incomplete trees (e.g., those significantly cut by bounding boxes), and incorrect detections. The final dataset consists of 2,900 TLS/ALS tree or tree-cluster pairs. Fig. 3 illustrates nine ALS/TLS example pairs, with ALS trees in yellow (top row) and corresponding TLS scans in aqua (bottom row). For validation, we construct a set that consists of 300 trees sampled from plots that are geographically distinct from those used for training and testing. These trees span the same range of mixed-conifer vegetation types but originate from separate ALS/TLS co-registered plots, ensuring that the model is evaluated on structurally similar but spatially independent examples. The validation set is used exclusively for hyperparameter tuning, monitoring reconstruction error during training, and selecting the best checkpoint under an early stopping criterion. Note that the test set is held out entirely until final evaluation, and no samples in the validation set appear in the training or test splits.
Figure 4: Empirical geometric relationship between ALS and TLS point clouds. ALS (in yellow) captures upper-canopy geometry, forming a sparse convex envelope (in green), while TLS points (in purple) representing stems, branches, and understory are largely contained within this envelope. This empirical property motivates the containment assumption exploited in this work when learning cross-domain mappings between ALS and TLS data.
@@
for a small constant $\epsilon\ll 1$, with $(\mathbf{X},\mathbf{Y})\sim\mathcal{D}$ drawn from the dataset of co-registered tree level TLS/ALS pairs. Here, $\mathbf{x}\sim\mathbf{X}$ are 3D points uniformly sampled from the TLS point cloud, and $1(\cdot)$ is the indicator function. In our empirical analysis of the training CoLIDAR-Forest3D (ALS/TLS) dataset, approximately 3.9% of TLS points fall outside the ALS convex hull, i.e., $\epsilon=0.04$, validating the assumption at the tree scale. We emphasize that this prior is not assumed to hold globally at the plot or landscape scale and that a variety of factors may violate containment. However, the small deviations from containment that may be locally present are expected and well justified, as they can result from sensor misalignment, resolution differences, vegetation visible from the ground but occluded from the aerial view, and inherent measurement noise in both ALS and TLS systems. Despite these localized discrepancies, the overall containment expectation introduces a geometric prior that is implicitly learned by our conditional generative model. This prior not only helps constrain synthesis within plausible ALS-guided structure, but also provides a useful means of measuring the spatial fidelity of the generated points relative to the ALS envelope.
@@
The DDPM architecture consists of two components: a forward diffusion process that progressively corrupts the input point cloud $\mathbf{X}^{0}\in\mathbb{R}^{N\times 3}$ over $T$ steps by multiplying it with a decaying factor while adding Gaussian noise, and a learned reverse process that denoises the corrupted sample back to a clean reconstruction. This yields a training trajectory: $\mathbf{X}^{0}\rightarrow\mathbf{X}^{1}\rightarrow\cdots\rightarrow\mathbf{X}^{T}\sim\mathcal{N}(0,\mathbf{I})$, as the forward (diffusion) process, $\mathbf{X}^{T}\rightarrow\mathbf{X}^{T-1}\rightarrow\cdots\rightarrow\hat{\mathbf{X}}^{0}$ as the reverse (denoising) process. The model conditions each denoising step on ALS observations $\mathbf{Y}$, enabling it to learn structure consistent with the canopy geometry while generating plausible TLS content beneath. The generative backbone used in ForestGen3D corresponds to the Conditional Generation Network (CGNet) architecture in [30], which enables us to condition denoising based on ALS point cloud measurements. CGNet consists of three principal components: a PointNet++ based denoising network [31], a ConditionNet that encodes ALS structure, and a FeatureMapper that fuses both representations. The denoising network follows a hierarchical, U-Net architecture that extracts features from noisy point clouds using multi-scale grouping and radius based neighborhood search. Four hierarchical levels are used, employing radii of 0.1, 0.2, 0.4, and 0.8 m with 32 neighbors sampled at each scale. Feature dimensions progress through $\{32,64,128,256,512\}$, and residual multi layer perceptron (MLP) blocks with skip connections are used to improve stability across diffusion steps. The diffusion timestep is embedded using a 128 dimensional positional encoding and concatenated to point level features. ALS-based conditioning is provided through the ConditionNet, which mirrors the hierarchical structure of the denoising network and extracts both local and global structural features from the ALS point cloud. Its feature dimensions follow $\{32,32,64,64,128\}$, enabling the network to capture the vertical distribution patterns, and scan sparsity. The resulting ALS features are aligned with the denoising hierarchy and supplied to the FeatureMapper, which integrates noisy point cloud features and ALS information using FiLM-style modulation [32]. This module applies shared MLP feature transforms, radius-based grouping, and an attention mechanism with grouped feature transformations and batch normalized activations. Together, these components ensure that the reverse diffusion process is guided both by ALS derived structure and by the learned statistical distribution of TLS geometry. A high-level schematic of the model architecture is shown in Fig. 5.
Figure 5: Schematic of ALS-to-TLS cross-domain ForestGen3D. The forward process encodes a TLS sample into Gaussian noise, while the reverse process learns to denoise a sample conditioned on ALS input following a denoising diffusion probabilistic model (DDPM) formulation, reconstructing high-resolution 3D vegetation structure.